OpenStudy (anonymous):

Calculus II - Volume of Solids - Setup Question \[y = 4x-x^2\] \[y=2-x\] Let D be the region bound by the curves above. Setup an integral representing the volume of the solid obtained by revolving R about each of the following lines; a) the x-axis b) the line y=5 c) the line y = -1 d) the line x = -1 e) the line x = 3 I have worked all five using the "washer method" and would like to get some feedback on the answers; a) \[\pi \int\limits_{-1}^{2} (4-x^2)^2 - (2-x)^2 \] b) \[\pi \int\limits_{-1}^{2} (2-x)^2 - (5-(4-x^2))^2 \] c) \[\pi \int\limits_{-1}^{2} (2+x)^2 - (1+(4-x^2))^2 \]

OpenStudy (anonymous):

d) \[\pi \int\limits_{-1}^{2} (2-x)^2 - (1+(4-x^2))^2 \] e) \[\pi \int\limits_{-1}^{2} (2-x)^2 - (3-(4-x^2))^2 \]

OpenStudy (amistre64):

dx's

OpenStudy (amistre64):

im assuming its the area bound by the intersection of these functions

OpenStudy (amistre64):

washer method you have to make sure your integrating and "adjusting" according to the axis of roataion as well

OpenStudy (amistre64):

a function viewed from the x axis doesnt look the same as a function viewed from the y axis; it becomes its inverse function

OpenStudy (anonymous):

Yes, bound by the intersection. Shoot, adjusting. I also tried this for D (shell method); d) \[2\pi \int\limits_{-1}^{2} (x+1)[(4-x^2)-(2-x)]\]

OpenStudy (anonymous):

So anytime I am rotating about the Y-axis I need to adjust my equations to be in terms of Y?

OpenStudy (anonymous):

Using the washer method, that is.

OpenStudy (amistre64):

adjust your equations according to axis and method of ...sheel, disc, washer etc

OpenStudy (amistre64):



OpenStudy (amistre64):

R = f(x) r = g(x) pi (R^2 - r^2) = washer area integrated with the x axis

OpenStudy (amistre64):

washer is just a double disc

OpenStudy (turingtest):

I meant shell*

OpenStudy (turingtest):

shell about x integrate w/respect to x disk about x integrate w/respect to y and vice-versa

OpenStudy (amistre64):

thats better :)

OpenStudy (amistre64):

and look into your function from the way the method is looking at it; if its a radius define it from the radial axis

OpenStudy (anonymous):

Ok, so if you rotate around y axis R = f(y) r = g(y)

OpenStudy (amistre64):

if its a heaight, define it from the base axis

OpenStudy (amistre64):

f-1(y) but yes

OpenStudy (amistre64):

inverse when you are looking at it from the other axis that the equation is defined for

OpenStudy (amistre64):

y=x^2 looking at it from the yaxis is x=sqrt(y)

OpenStudy (amistre64):